Dr. Grey wanted to see if doctors prescribed different treatments for OCD for male patients compared to females. Please use the following frequency tables for the data collected by Dr. Grey.

Contingency table of Observed and Expected Frequencies of the different treatments (Drug vs. Cognitive Therapy) prescribed for patients with OCD.

Therapy Totals

Drug

Cognitive Therapy

Total

Gender of Patient

Observed

Expected

Observed

Expected

Man

20

10

30

Woman

15

15

30

Total

35

25

60

What is the critical value?

A) 3.84

B) 5.99

C) 4.00

D) 2.71
Answer #1

This test is independent chi square test.

Here we do not have level of significance, so default it is 0.05.

Therapy Totals Drug Cognitive Therapy
Gender of patient Man 20 10
Woman 15 15

We have 3 rows and 2 columns.

So df = ( r-1)(c-1) = ( 2-1)(2-1) = 1*1 = 1

Hence the critical value is,

-----------( Using excel formula " =chisq.inv.rt(0.05,1)" )

Hence the correct option is A) 3.84.